# mO [#1]Paola

### v1

@@ Abstract
We introduce AutoVER, an Autoregressive model for Visual Entity Recognition. Our model extends an autoregressive Multimodal Large Language Model by employing retrieval augmented constrained generation. It mitigates low performance on out-of-domain entities while excelling in queries that require visually-situated reasoning. Our method learns to distinguish similar entities within a vast label space by contrastively training on hard negative pairs in parallel with a sequence-to-sequence objective without an external retriever. During inference, a list of retrieved candidate answers explicitly guides language generation by removing invalid decoding paths. The proposed method achieves significant improvements across different dataset splits in the recently proposed Oven-Wiki benchmark. Accuracy on the Entity seen split rises from 32.7% to 61.5%. It also demonstrates superior performance on the unseen and query splits by a substantial double-digit margin.
@@ 1 Introduction
Multimodal Large Language Models (Multimodal LLMs) have demonstrated superior performance on a variety of vision-and-language tasks such as visual question answering, image captioning, zero-shot image classification, among others [46, 33]. Their abilities can be transferred with few-shot tuning [59, 2] or even learning in context without parameter updates [4, 62]. Given their remarkable generalization abilities and prior knowledge based on large scale pretraining we consider there is still great potential for using them on tasks that require knowledge grounding.
A recently proposed task requiring knowledge grounding is the Open-domain Visual Entity Recognition (Oven-Wiki) [23] task. In this task, given an input image and a question about the image, the goal is to answer with a very specific entity from Wikipedia. For instance, the answers could be a specific model of airplane, e.g. ATR 42, or British Aerospace 146. This is a very challenging task where Multimodal LLM-based solutions are prone to hallucinations, or producing answers that are not at the right level of granularity. Importantly, visual entity recognition requires recognizing entities that never appear in the training data (unseen split). Moreover, a portion of the Oven-Wiki benchmark (query split) extends the task beyond recognition, requiring non-trivial reasoning to resolve the query question.
@@
Visual Entity Recognition (VER) presents unique challenges compared to general Visual Question Answering (VQA) [1]. The first challenge lies in the answer label space comprising over 6 million Wikipedia entities, which is prohibitive for classifier-based VQA models [49] as many entities considered in this ontology are visually very similar. The second issue involves generation-based VQA solutions [33, 14], where hallucinations can lead to generated text that is not grounded to the entity space. Lastly, existing VQA approaches fail to consider the visual information of candidates [56, 35]. Entity images from the knowledge base also play a significant role when disambiguating between entities with similar identifiers but different visual appearances. These methods also struggle with out-of-domain generalization and multi-hop reasoning, challenges which are precisely covered by the unseen and query split of the Oven-Wiki dataset.
In this work, we introduce an Autoregressive Visual Entity Recognizer, the first approach that enables multimodal language models to perform accurate visual entity recognition over a massive knowledge base, abbreviated as AutoVER. AutoVER addresses entity recognition by reformulating the problem as a sequence-to-sequence generation problem, as depicted in Fig. 1. The query image is translated into the token embedding space using a learnable projection layer, in a manner akin to treating images as a foreign language [37, 61]. This allows us to utilize pre-trained multi-modal language models and enormously improves performance on the query split which requires reasoning.
@@
In summary, AutoVER offers several key advantages over baselines and prior work on VQA in the realm of visual entity recognition: (i) we refine the recognition process by integrating contrastive training into an MLLM. (ii) our proposed retrieval-augmented constrained decoding framework guarantees correct grounded entity prediction and enhances prediction over unseen entities; (iii) we show that leveraging pre-trained visual models and knowledge graphs for hard negative mining significantly strengthens our contrastive-generative framework in fine-grained entity recognition; (iv) experimental results show that AutoVER consistently outperforms fine-tuned CLIP and PaLI variants on all Oven-Wiki splits. For instance, AutoVER-7B achieves a 61.5% accuracy on the Oven-Wiki entity seen split over PaLI-17B’s 30.6%, and 21.7% on entity unseen over PaLI-17B’s 12.4%. Additionally, we evaluate AutoVER alongside public multimodal LLMs on entity-related questions from the A-OKVQA, demonstrating that our model can effectively zero-shot transfer to general out-of-domain VQA datasets. Furthermore, ablation studies confirm the effectiveness of the retrieve-generate framework we introduce for visual entity recognition. Code will be released at https://github.com/MrZilinXiao/AutoVER.
@@ 2 Related Work
Visual Entity Recognition (VER) [23] is an emerging task for assessing the ability of a model to perform multi-modal knowledge grounding. It can be regarded as a multi-choice variant of Visual Question Answering (VQA) [1] with the key distinction that the choice set will comprise all entities in a specific knowledge base. Earlier solutions view this problem as an image-text-to-image-text retrieval task, where CLIP-based models [47] are fine-tuned and the top-scored answer in inference is treated as the model prediction. Alternatively, others rely on fine-tuning generative language models such as PaLI [9] in an attempt to match the generated text with candidate entity identifiers using sparse matching approaches such as BM25 [52]. Entity Linking (EL) has been a long-standing text-only counterpart to VER, which entails locating mentions in the document and disambiguating them against a set of candidate entities. Similar to prior solutions in VER, a bi-encoder was used to retrieve and link each potential mention to an external knowledge base. Our method bears the closest resemblance with the recent generative EL paradigm [6, 40, 67, 68] which also reduces the knowledge-grounding problem into a sequence-to-sequence generation task. However, we stand out with several distinct advantages, including the seamless integration of retrieval capabilities into the language model and the usage of explicit guidance for sequence generation.
@@
Multi-modal Large Language Models (MLLMs) are motivated by the remarkable reasoning abilities of Large Language Models (LLMs), enabling researchers to explore transferring them into the vision and other domains. BLIP [34] proposes to bootstrap the multi-modal understanding ability into a unified semantic space. Flamingo [2] applies a cross-attention architecture to fuse visual contexts, thus bringing visual in-context learning. LLaVA [37] builds upon LLaMA [58] and uses instruction-tuning to align visual features to the language space. The common practice of prompting MLLMs to produce proxy representations for downstream usage is to expand the vocabulary of language models. Such expansion augments the MLLM functionality beyond language generation. FROMAGe [27], GILL [26], and LISA [30] achieve satisfactory results over image-text-to-text retrieval, image generation and reasoning segmentation by inserting a retrieval token, a generation token and a segmentation token into MLLMs respectively. Our approach, built upon the latest MLLM model architecture, presents a compelling alternative to the bi-encoder shallow dot-product interaction or encoder-decoder sparse surface form matching employed in previous works on visual entity recognition.
@@
Retrieval-Augmented Language Models (RALMs) represent a class of NLP solutions focusing on knowledge-intensive tasks. RALMs typically condition a language model on relevant documents from a grounding corpus during generation, thereby enhancing the performance in knowledge-intensive language understanding tasks. Lewis et al. [32] jointly fine-tune a retriever with an encoder-decoder model, enabling the community to explore the RALM paradigm in language understanding. Guu et al. [19] train a bi-directional variant and also demonstrate superior performance. Ram et al. [48] propose to do RALM in-context without any further training of the LM. Apart from retrieved items in context, Févry et al. [13] are the first to integrate retrieved entity supervision by injecting intermediate representations. While augmented context helps in mitigating the well-known issue of hallucination, it does not ensure a faithful prediction with respect to an external knowledge base. Our retrieve-generate framework is distinct from existing retrieval-augmented methods. AutoVER reduces the candidate entity set from millions to hundreds for generative language models through retrieval and dynamically constrains language model generation using a prefix tree. This framework coincides with the design philosophy of agents [18] in Interactive NLP [63], where an agent interacts with the dynamic environment by beam-searching all available options.
@@ 3.1 Problem Definition
Open-domain Visual Entity Recognition (VER) can be viewed as a multimodal knowledge grounding task, in which the model is required to process an image-text pair input $x=(\mathbf{Q}_{\text{im}},\mathbf{Q}_{\text{t}})$ and predict an entity $e$. $\mathbf{Q}_{\text{t}}$ describes the specific intent that prompts the model to ground some entity $e$ in the image $\mathbf{Q}_{\text{im}}$ to a label space $\mathcal{E}$. Each entity $e\in\mathcal{E}$ is a member of the knowledge base $\mathcal{K}=\{(e,\mathbf{E}_{\text{im}},\mathbf{E}_{\text{desc}})\mid e\in\mathcal{E}\}$ where $\mathbf{E}_{\text{desc}}$ are the text description and $\mathbf{E}_{\text{im}}$ is a set of relevant images of the entity. To circumvent trivial solutions for some query questions, all image-text pairs are annotated so the question cannot be correctly answered without the image.
Considering that Oven-Wiki is the only dataset currently available for VER, we highlight the dataset splits within Oven-Wiki to underscore the challenges of the VER task. The dataset is divided into train, validation, test, and human-validation sets, with each of the first three containing both an entity split and a query split. The Entity split focuses on image recognition across diverse visual entities, while the Query split pertains to Visual Question Answering with visually-situated natural language queries. Each split is further divided into seen and unseen subsets to assess the generalization capabilities of models beyond the training domain.
@@ 3.2 Model Overview
$\mathbf{H}_{\text{im}}=\mathbf{W}_{q}\cdot g_{\text{vision}}\left(\mathbf{Q}_{\text{im}}\right)$
$\mathbf{H}_{\text{t}}=\mathbf{W}_{\text{embed}}\cdot\mathbf{Q}_{\text{t}}$
@@
On the entity encoder side, a frozen CLIP visual encoder $g_{\text{vision}}$ transforms entity image $\mathbf{E}_{\text{im}}$ into grid image features $\mathbf{Z}_{\text{im}}$, and the frozen CLIP text encoder $g_{\text{text}}$ encodes the entity identifier and description into text features $\mathbf{Z}_{\text{text}}$. A two-layer Transformer encoder handles the fusion of two modalities by a fusion token as a soft prompt [31], whose design was also adopted in [25]. The final hidden states of the fusion token are also subjected to dimension matching and L2 normalization, providing the entity representation $E$.
@@
In contrast to image-text retrieval which offers a unique correspondence between images and texts, query-to-entity mapping is subjective, i.e. multiple queries in the training data are associated with the same entity. This requires a particular sampling strategy during training by ensuring that a batch does not include the same entity, referred to as a “conflicting batch”. To preserve the integrity of the training data distribution, we implement a rejection sampling [16] method as an alternative to standard random sampling, resorting to resampling upon encountering conflicting batches.
@@
Figure 3: Retrieval-augmented constrained decoding illustration of AutoVER inference process. The representation $Q$ will query a pre-cached entity database constructed using the multimodal entity encoder, and get the top-$k$ candidate entities. A prefix-constrained tree is dynamically built based on retrieved entity identifiers and guides the language model to autoregressively generate the next token, thereby ensuring the successful grounding of generated content.
@@ 3.3 Hard-negative Mining
To alleviate the entity granularity problem, two hard negative sampling approaches are proposed in our contrastive-generative framework, i.e. vision-hard and kb-hard. Both approaches create entity groups used for constructing an in-batch sampler that prefers sampling similar entities in contrastive training, yet they differ in how they generate these similar groups. vision-hard relies on a pre-trained ViT [12] image classifier to identify visually similar entities based on shared prediction classes. In contrast, kb-hard uses the category hierarchy of Wikidata as an external knowledge source, considering entities that share a parent node in the category hierarchy as knowledge-similar entities. We refer readers to Appendix 0.B for details about the selection of hard negative groups.
@@ 3.4 Retrieval-augmented Constrained Decoding
We illustrate the model inference in Fig. 3. After training, all entity candidates are cached using the multi-modal entity encoder $F_{\varphi}$ to construct an entity vector database $\mathcal{V}\in\mathbb{R}^{n\times d}$ for efficient retrieval. Given an evaluation sample $\left(\mathbf{Q}_{\text{im}},\mathbf{Q}_{\text{t}}\right)$, we first forward the MLLM for the normalized representation of the <ret> token, and query database $\mathcal{V}$ using top-$k$ similarity search to get $k$ entity candidates. The model dynamically generates a prefix tree (trie) that covers $k$ entity candidates, and the trie will explicitly guide entity identifier generation by eliminating impossible decoding paths when autoregressively generating tokens. The retrieval process takes into consideration the image of entities and improves the model towards out-of-domain entities. Moreover, the retrieval-augmented constrained decoding guarantees the grounding of the generated content to the knowledge base, alleviating the issue of hallucinations.
@@ 4.1 Settings
Metrics. We follow the standard setup in [23], which uses accuracy and harmonic mean to evaluate model performance on different data splits. The harmonic mean of each split will equally weigh the importance of the seen and unseen subsets and penalize models that show a weakness in either aspect. Finally, we report the overall harmonic mean on the Entity and Query splits as the final metric for the validation and test sets.
Data Pre-processing and Models. We pre-process all query and entity images by resizing them into 336 $\times$ 336 pixels with padding to keep an identical aspect ratio. Entity descriptions are truncated to 77 tokens to fit the context window size of CLIP by cutting off sentences to the maximum available ones. Since query texts are typically short, no truncation is needed for query texts. Encoders on the entity side and visual encoder in MLLM are pre-trained CLIP-ViTL/14-336px. MLLM is initialized with the vicuna-7b-v1.5 or vicuna-13b-v1.5 checkpoint and corresponding visual projectors pre-aligned on a 558k subset of LAION [54], CC [7] and SBU [45] curated by [36].
Training and Evaluation. We train on the entity train and query train splits of Oven-Wiki, which consists of nearly 5 million query-entity pairs. We refer to Appendix 0.A for hyper-parameter choices and training details. We set $\lambda_{r}$ to 1 for all training settings. Due to a limited compute budget, we select 10% and 50% of the training data through weighted sampling according to entity occurrence frequency to facilitate ablation studies and final results reporting. For the same reason, we do not run a hyperparameter search. In evaluation, the number of retrieved entities $k$ is set to 300 according to empirical trials on the validation split.

### v2

@@ Abstract
We introduce AutoVER, an Autoregressive model for Visual Entity Recognition. Our model extends an autoregressive Multimodal Large Language Model by employing retrieval augmented constrained generation. It mitigates low performance on out-of-domain entities while excelling in queries that require visual reasoning. Our method learns to distinguish similar entities within a vast label space by contrastively training on hard negative pairs in parallel with a sequence-to-sequence objective without an external retriever. During inference, a list of retrieved candidate answers explicitly guides language generation by removing invalid decoding paths. The proposed method achieves significant improvements across different dataset splits in the recently proposed Oven-Wiki benchmark with accuracy on the Entity seen split rising from 32.7% to 61.5%. It demonstrates superior performance on the unseen and query splits by a substantial double-digit margin, while also preserving the ability to effectively transfer to other generic visual question answering benchmarks without further training.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have demonstrated superior performance on a variety of vision-and-language tasks such as visual question answering, image captioning, zero-shot image classification, among others [55, 40]. Their abilities can be transferred with few-shot tuning [69, 2] or even learning in context without parameter updates [5, 73]. Given their remarkable generalization abilities and prior knowledge based on large-scale pre-training, we consider there is still great potential for using them on tasks that require knowledge grounding.
A recently proposed task requiring knowledge grounding is the Open-domain Visual Entity Recognition (Oven-Wiki) [28] task. In this task, given an input image and a question about the image, the goal is to answer with a very specific entity from Wikipedia. For instance, the answers could be a specific model of airplane, e.g. ATR 42, or British Aerospace 146. This is a very challenging task where MLLM-based solutions are prone to hallucinations, or producing answers that are not at the right level of granularity. Importantly, visual entity recognition requires recognizing entities that never appear in the training data (unseen split). Moreover, a portion of the Oven-Wiki benchmark (query split) extends the task beyond recognition, requiring non-trivial reasoning to resolve the query question.
@@
Visual Entity Recognition (VER) presents unique challenges compared to general Visual Question Answering (VQA) [45, 1]. The first challenge lies in the answer label space comprising over 6 million Wikipedia entities, which is prohibitive for classifier-based VQA models [58] as many entities considered in this ontology are visually very similar. The second issue involves generation-based VQA solutions [40, 18], where hallucinations can lead to generated text that is not grounded to the entity space. Lastly, existing VQA approaches fail to consider the visual information of candidates [65, 41]. Entity images from the knowledge base also play a significant role when disambiguating between entities with similar identifiers but different visual appearance. These methods also struggle with out-of-domain generalization and multi-hop reasoning, challenges which are covered by the unseen and query split of the Oven-Wiki dataset.
In this work, we introduce an Autoregressive Visual Entity Recognizer (AutoVER), the first approach that enables multimodal language models to perform accurate visual entity recognition over a massive knowledge base. AutoVER addresses entity recognition by reformulating the problem as a sequence-to-sequence generation problem, as depicted in Fig. 1. The query image is translated into the token embedding space using a learnable projection layer, in a manner akin to treating images as a foreign language [43, 72]. This allows us to utilize pre-trained multi-modal language models and enormously improves performance on the query split of Oven-Wiki which requires reasoning over spatial relations, commonsense, and other visually-situated contexts.
@@
In summary, AutoVER offers several advantages over baselines and prior work on VQA in the realm of visual entity recognition: (i) It refines the recognition process by integrating contrastive training into an MLLM. (ii) The proposed retrieval-augmented constrained decoding framework guarantees correct grounded entity prediction and enhances prediction over unseen entities; (iii) AutoVER leverages pre-trained visual models and knowledge graphs for hard negative mining which significantly strengthens our contrastive-generative framework in fine-grained entity recognition; (iv) Experimental results show that AutoVER consistently outperforms fine-tuned CLIP and PaLI variants on all Oven-Wiki splits. For instance, AutoVER-7B achieves a 61.5% accuracy on the Oven-Wiki entity seen split over PaLI-17B’s 30.6%, and 21.7% on entity unseen over PaLI-17B’s 12.4%. Additionally, we evaluate AutoVER alongside public multimodal LLMs on entity-related questions from the A-OKVQA dataset, demonstrating that our model can effectively zero-shot transfer to out-of-domain VQA datasets beyond Oven-Wiki. Furthermore, ablation studies confirm the effectiveness of the introduced retrieve-generate framework. Code is released at https://github.com/MrZilinXiao/AutoVER.
@@ 2 Related Work
Visual Entity Recognition (VER) [28] is an emerging task for assessing the ability of a model to perform multi-modal knowledge grounding [29, 80, 53]. It can be regarded as a variant of visual question answering [45, 59, 1] with the key distinction that the choice set will comprise all entities in a specific knowledge base. Earlier solutions view this problem as an image-text-to-image-text retrieval task, where CLIP-based models [56] are fine-tuned and the top-scored answer in inference is treated as the model prediction. Alternatively, others rely on fine-tuning generative language models such as PaLI [12] and GiT [71, 8] in an attempt to match the generated text with candidate entity identifiers using sparse matching approaches such as BM25 [61]. Entity Linking (EL) has been a long-standing text-only counterpart to VER, which entails locating mentions in the document and disambiguating them against a set of candidate entities. Our method bears the closest resemblance with the recent generative EL paradigm [7, 47, 78, 79], which also reduces the knowledge-grounding problem into a sequence-to-sequence generation task. However, we stand out with several distinct advantages, including the seamless integration of retrieval capabilities into the language model and the use of explicit guidance for sequence generation.
@@
Multimodal LLMs (MLLMs) are motivated by the remarkable reasoning abilities of Large Language Models (LLMs). LLaVA [43] builds upon LLaMA [68] and uses instruction-tuning to align visual features to the language space. The common practice of prompting MLLMs to produce proxy representations for downstream usage is to expand the vocabulary of language models. Such expansion augments the MLLM functionality beyond language generation. FROMAGe [34], GILL [33], GenLLaVa [26] and LISA [37] achieve satisfactory results over image-text-to-text retrieval, image generation and reasoning segmentation by inserting special tokens to augment the MLLMs original functionality. Our approach, built upon one of the latest MLLM model architecture, presents a compelling alternative to the bi-encoder shallow dot-product interaction or encoder-decoder sparse surface form matching employed in previous works on visual entity recognition. Our method also integratives contrastive learning on the outputs of entities and question pairs, analogous to the image-text contrastive learning used in CLIP-like models [56, 66, 54] and among visual features used for self-supervised contrastive learning [11, 25, 3, 16].
@@
Retrieval-Augmented Language Models (RALMs) represent a class of NLP solutions focusing on knowledge-intensive tasks. RALMs typically condition a language model on relevant documents from a grounding corpus during generation, thereby enhancing the performance in knowledge-intensive language understanding tasks. Lewis et al. [39] jointly fine-tune a retriever with an encoder-decoder model, enabling the community to explore the RALM paradigm in language understanding. Guu et al. [23] train a bi-directional variant and also demonstrate superior performance. Apart from retrieved items in context, Févry et al. [17] are the first to integrate retrieved entity supervision by injecting intermediate representations. While augmented context helps in mitigating the well-known issue of hallucination, it does not ensure a faithful prediction with respect to an external knowledge base. Our retrieve-generate framework is distinct from existing retrieval-augmented methods. AutoVER reduces the candidate entity set from millions to hundreds for generative language models through retrieval [30] and dynamically constrains language model generation using a prefix tree. This framework coincides with the design philosophy of agents [22] in Interactive NLP [74], where an agent interacts with the dynamic environment by performing beam-search on all available options.
@@ 3 Methodology
We first formulate the Visual Entity Recognition problem in Sec. 3.1. Then we present the overview design of AutoVER in Sec. 3.2. We illustrate hard-negative mining with knowledge base source and pre-trained visual model in Sec. 3.3 and retrieval-augmented constrained decoding process in Sec. 3.4.
@@ 3.1 Problem Definition
Visual Entity Recognition can be viewed as a multimodal knowledge grounding task, in which the model is required to process an image-text pair input $x=(\mathbf{Q}_{\text{im}},\mathbf{Q}_{\text{t}})$ and predict an entity $e$. $\mathbf{Q}_{\text{t}}$ describes the specific intent that prompts the model to ground some entity $e$ in the image $\mathbf{Q}_{\text{im}}$ to a label space $\mathcal{E}$. Each entity $e\in\mathcal{E}$ is a member of the knowledge base $\mathcal{K}=\{(e,\mathbf{E}_{\text{im}},\mathbf{E}_{\text{desc}})\mid e\in\mathcal{E}\}$ where $\mathbf{E}_{\text{desc}}$ is a text description and $\mathbf{E}_{\text{im}}$ is a set of relevant images of the entity. To circumvent trivial solutions for some query questions, all image-text pairs are annotated so the question cannot be correctly answered without the image.
@@ 3.2 Model Overview
$\mathbf{H}_{\text{im}}=\mathbf{W}_{q}\cdot g_{\text{vision}}\left(\mathbf{Q}_{\text{im}}\right),$
$\mathbf{H}_{\text{t}}=\mathbf{W}_{\text{embed}}\cdot\mathbf{Q}_{\text{t}},$
@@
On the entity encoder side, a frozen CLIP visual encoder $g_{\text{vision}}$ transforms entity image $\mathbf{E}_{\text{im}}$ into grid image features $\mathbf{Z}_{\text{im}}$, and the frozen CLIP text encoder $g_{\text{text}}$ encodes the entity identifier and description into text features $\mathbf{Z}_{\text{text}}$. A two-layer Transformer encoder handles the fusion of two modalities by a fusion token as a soft prompt [38], whose design was also adopted in [32]. The final-layer hidden states of the fusion token are also subjected to dimension matching and L2 normalization, providing the entity representation $E$.
@@
In contrast to image-text retrieval which offers a unique correspondence between images and texts, query-to-entity mapping is subjective, i.e. multiple queries in the training data are associated with the same entity. This requires a particular sampling strategy during training to ensure that a batch does not include the same entity, referred to as a “conflicting batch”. To preserve the integrity of the training data distribution, we implement a rejection sampling [20] method as an alternative to standard random sampling, resorting to resampling upon encountering conflicting batches.
@@
Figure 3: Retrieval-augmented constrained decoding illustration of our proposed AutoVER inference process. The representation $Q$ will query a pre-cached entity database constructed using the multimodal entity encoder, and get the top-$k$ candidate entities. A prefix-constrained tree is dynamically built based on retrieved entity identifiers and guides the language model to autoregressively generate the next token, thereby ensuring the successful grounding of generated content.
@@ 3.3 Hard-Negative Mining
To alleviate the entity granularity problem, two hard negative sampling approaches are proposed in our contrastive-generative framework, i.e. vision-hard and kb-hard. Both approaches create entity groups used for constructing an in-batch sampler that prefers sampling similar entities in contrastive training, yet they differ in how they generate these similar groups. vision-hard relies on a pre-trained ViT [15] image classifier to identify visually similar entities based on shared prediction classes. In contrast, kb-hard uses the category hierarchy of Wikidata as an external knowledge source, considering entities that share a parent node in the category hierarchy as knowledge-similar entities. We refer readers to the supplementary material for details about the construction of hard negative groups.
@@ 3.4 Retrieval-augmented Constrained Decoding
We illustrate the proposed model inference in Fig. 3. After training, all entity candidates are cached using the multi-modal entity encoder $F_{\varphi}$ to construct an entity vector database $\mathcal{V}\in\mathbb{R}^{n\times d}$ for efficient retrieval. Given an evaluation sample $\left(\mathbf{Q}_{\text{im}},\mathbf{Q}_{\text{t}}\right)$, we first forward the MLLM for the normalized representation of the <ret> token, and query database $\mathcal{V}$ using top-$k$ similarity search to get $k$ entity candidates. The model dynamically generates a prefix tree (trie) that covers $k$ entity candidates, and the trie will explicitly guide entity identifier generation by eliminating impossible decoding paths when autoregressively generating tokens. The retrieval process takes into consideration the image of entities and improves the model towards out-of-domain entities. Moreover, the retrieval-augmented constrained decoding guarantees the grounding of the generated content to the knowledge base, alleviating the issue of hallucinations.
@@ 4 Experiments
We describe the experimental setting in Sec. 4.1 and report the main results in Sec. 4.2. We present the zero-shot generalization results in Sec. 4.3 and ablation study in Sec. 4.4 with discussions. For an intuitive demonstration of our model, we present the case study in Sec. 4.5.
@@ 4.1 Settings
Metrics. We follow the standard setup in Hu et al. [28], which uses accuracy and harmonic mean to evaluate model performance on different data splits. The harmonic mean of each split will equally weigh the importance of the seen and unseen subsets and penalize models that show weakness in either aspect. Finally, we report the overall harmonic mean on the Entity and Query splits as the final metric for the validation and test sets.
Data Pre-processing and Models. We pre-process all query and entity images by resizing them into 336 $\times$ 336 pixels with padding to keep an identical aspect ratio. Entity descriptions are truncated to 77 tokens to fit the context window size of CLIP by cutting off sentences to the maximum available ones. Since query texts are typically short, no truncation is needed for query texts. Encoders on the entity side and visual encoder in MLLM are pre-trained CLIP-ViTL/14-336px. The MLLM is initialized with the vicuna-7b-v1.5 or vicuna-13b-v1.5 checkpoint and corresponding visual projectors pre-aligned on a 558k subset of LAION [63], CC [9] and SBU [52] curated by Liu et al. [42].
Training and Evaluation. We train on the entity train and query train splits of Oven-Wiki, which consists of nearly 5 million query-entity pairs. We conduct all pieces of training in a batch size of 256 on 32 V100-SXM2-32GB GPUs. We refer to the supplementary material for hyper-parameter choices and training details. We set $\lambda_{r}$ to 1 for all training settings. Due to a limited compute budget, we select 10% and 50% of the training data through weighted sampling according to entity occurrence frequency to facilitate ablation studies and reporting final results. For the same reason, we do not run a hyperparameter search. In evaluation, the number of retrieved entities $k$ is set to 300 according to empirical trials on the validation split.